OpenStudy (anonymous):

solve for "k" 8/7k+8/9=3- -3/5k can someone explain the steps i have to take

OpenStudy (anonymous):

\[\frac{8}{7}k + \frac{8}{9}= 3 - - \frac{3}{5}k\] is it this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first move k onto one side

OpenStudy (anonymous):

i get \[\frac{ 39 }{ 35 }k+\frac{ 8 }{ 9 }=3\]

OpenStudy (anonymous):

because there was 2 negative sign - - before 3k/5 so it's a +. \[\frac{8}{7}k-\frac{3}{5}k+\frac{8}{9}=3\]\[\frac{40}{35}k-\frac{21}{35}k+\frac{8}{9}=3\]\[\frac{19}{35}k + \frac{8}{9} = 3\]

OpenStudy (anonymous):

oh ok so from there i subtract \[\frac{ 8 }{ 9 }\] from 3 and i got 11/9

OpenStudy (anonymous):

or is it 10/9

OpenStudy (anonymous):

\[3-\frac{8}{9} = \frac{27}{9} - \frac{8}{9} = \frac{19}{9}\] \[\frac{19}{35}k = \frac{19}{9}\]

OpenStudy (anonymous):

so from there i have to divide \[\frac{ 19 }{ 35 }\] from both sides

OpenStudy (anonymous):

I get for the answer i get 665/171

OpenStudy (anonymous):

reduced its 35/9

OpenStudy (anonymous):

thank you very much i really appreciate it

OpenStudy (anonymous):

yup ah noproblems